OpenStudy (anonymous):

Express f(x) in the form a(x − h)2 + k. f(x) = − 3/4x^2 + 15x − 78

OpenStudy (raden):

f(x) = − 3/4x^2 + 15x − 78 f(x) = -3/4 (x^2 - 20x + 104) f(x) = -3/4 (x^2 - 20x + 100 + 4) f(x) = -3/4 (x^2 - 20x + 100) - (3/4)*(4) f(x) = -3/4 (x-10)^2 - 3

OpenStudy (anonymous):

Thanks :)

OpenStudy (raden):

welcome